# Hyperbolic embedding of brain networks as a tool for epileptic seizures forecasting

### v3

@@ I Introduction
Existing methods for predicting epileptic seizures primarily depend on long-term electroencephalographic (EEG) recordings, whether from the scalp or intracranial sources. This approach places considerable demands on both data collection and processing [12, 13]. These models are designed to identify specific preictal changes associated with each seizure, without accounting for alterations caused by shifts in vigilance states. However, such methods necessitate either real-time detection of vigilance states or the integration of various reference interictal states. While some methods achieve high sensitivity, they are often associated with high false positive rates, resulting in frequent false alarms that can diminish their clinical utility [14, 15]. Additionally, performance varies significantly across studies due to differences in datasets, feature selection, and model architectures, underscoring the absence of a standardized, generalizable solution for seizure prediction. Despite these challenges, recent advancements in brain network analysis have created new opportunities for developing more efficient and interpretable predictive models.
Previous studies [16, 17] have shown that brain connectivity matrices derived from intracranial EEG data can effectively identify preictal states. Building on these findings, we introduce a novel seizure forecasting approach that addresses key limitations of existing methods. Unlike continuous monitoring systems, our approach focuses on two main objectives: (1) determining whether patient-specific EEG connectivity during vigilance-controlled resting-state periods differs between seizure and non-seizure days, and (2) prospectively evaluating its relevance for daily seizure risk forecasts using calibrated predictive models. To estimate seizure risk, we utilize a hyperbolic network embedding method to represent patients’ functional connectivity in a low-dimensional space. This representation leverages hierarchical, small-world, and other key properties of brain networks, offering a more suitable framework for capturing connectivity patterns associated with seizure occurrence. Our method outperforms various forecasting models, including traditional machine learning approaches. By identifying robust biomarkers for seizure forecasting, our approach requires less training data while delivering superior performance, paving the way for practical and efficient solutions in epilepsy management.
@@ III.1 Discrimination of preictal networks
We first evaluated the ability of the proposed method to discriminate between interictal and preictal networks. Classification performances were assessed by AUC, the accuracy, and the F1-score and are presented in Tab. 1. In general, the connectivity patterns associated with activities of separated frequency bands cannot distinguish between preictal and interictal days. Only networks estimated from low $\gamma$ oscillations (30-49 Hz) performed better than a random classifier. Nevertheless, classification performances considerably increased when the information from different frequency bands was combined for each patient: All patients reached AUCs of $\geqslant 0.78$ (mean AUC $=0.93$, mean F1 score $=0.79$, and mean accuracy $=0.85$). These findings suggest that combining connectivity across different frequency bands provides valuable information for distinguishing preictal from interictal epochs. Notably, the hyperbolic embedding method outperformed a standard machine learning model, specifically a support vector machine (SVM). The SVM model achieved an average AUC of $0.79$, an F1 score of $0.78$, and a mean accuracy of $0.77$, highlighting the superior performance of the hyperbolic embedding approach.
@@ III.2 Forecasting of upcoming seizure(s)
Forecasting results were consistent with those of our previous study, in which we demonstrated that brain connectivity obtained from vigilance-controlled resting-state recordings could accurately predict the risk of upcoming seizures. [18]. The results, presented in Table. 2, demonstrate the forecasting performance obtained for the distinct frequency bands when the prediction models for each day were trained with data from all preceding days. Combining information from different frequency bands yielded a good predictor with an average Brier score and BSS of $0.12$ [min. = 0.003, max. = 0.48] and $0.36$ [min. $=-0.43$, max. $=0.70$], respectively. In comparison, the SVM model achieved a mean Brier score of $0.13$. It is worthy of notice that connectivity networks from the $\beta$ and low $\gamma$ band also provided good predictors, in agreement with previous findings [18]. These results demonstrate that our prediction model outperformed other approaches, including a standard machine learning model, in forecasting the risk of upcoming seizure(s).
@@ IV Discussion
Unlike current models that rely on long-term EEG recordings to track changes preceding seizures, our findings indicate that short resting-state recordings can offer valuable insights into the preictal state on a daily scale. It is worth noting that a direct comparison to traditional seizure prediction methods is not entirely feasible due to methodological differences, particularly in the duration of recordings utilized. However, compared to alternative models such as the SVM classifier [18] or a simple anteriority-based model, our approach exhibits superior performance. Additionally, most EEG-based methods for seizure prediction provide forecasts within minutes to, at most, an hour before a seizure event [14, 15]. In contrast, our method provides a single daily forecast, offering an extended time window for intervention or monitoring throughout the day.
@@
Despite the good classification and prediction results, our study has some limitations. First, the vigilance-controlled resting-state iEEG recordings were obtained from hospitalized patients who were candidates for presurgical evaluation. Although no trivial correlations were found between the occurrence of seizures and the controlled medication tapering [18], the results cannot be directly generalized to out-of-hospital real-life conditions. Similarly, given the heterogeneity of electrode location among the patients, we cannot generalize a single unique prediction model. While a larger cohort of patients is necessary to map out the limits of the proposed approach, the results do suggest that the method has the potential to be a promising tool for the analysis of brain networks.
To summarize, our findings indicate that non-Euclidean embeddings of brain networks may offer more discriminative information for distinguishing between interictal and preictal states, thereby providing a promising tool for accurately forecasting the daily risk of upcoming seizure(s).

### v4

@@ I Introduction
Current methods for predicting epileptic seizures largely rely on long-term electroencephalographic (EEG) recordings, obtained either from scalp or intracranial sources. These techniques impose significant demands on both data collection and processing [12, 13]. These models are designed to identify specific preictal changes associated with each seizure, without considering fluctuations caused by shifts in vigilance states. As a result, they require either real-time monitoring of vigilance levels or the incorporation of diverse interictal reference states While some methods achieve high sensitivity, they are often associated with high false positive rates, resulting in frequent false alarms that can diminish their clinical utility [14, 15]. Additionally, performance varies significantly across studies due to differences in datasets, feature selection, and model architectures, underscoring the absence of a standardized, generalizable solution for seizure prediction. Despite these challenges, recent advancements in brain network analysis have created new opportunities for developing more efficient and interpretable predictive models.
Previous studies [16, 17] have shown that brain connectivity matrices derived from intracranial EEG data can effectively identify preictal states. Building on these findings, we introduce a novel seizure forecasting approach that addresses key limitations of existing methods. In contrast to continuous monitoring systems, our approach is centered on two primary objectives: (1) assessing whether patient-specific EEG connectivity during vigilance-controlled resting-state periods differs between seizure and non-seizure days, and (2) prospectively evaluating the utility of these differences for forecasting daily seizure risk through calibrated predictive models. To estimate seizure risk, we utilize a hyperbolic network embedding method to represent patients’ functional connectivity into a low-dimensional space. This representation capture subtle changes in brain networks associated with seizure occurrence. Our method demonstrates superior performance compared to various forecasting models, including conventional machine learning techniques. By identifying robust biomarkers for seizure prediction, our approach requires less training data while achieving enhanced accuracy, offering a practical and efficient pathway toward improved epilepsy management.
@@ III.1 Discrimination of preictal networks
We first evaluated the ability of the proposed method to discriminate between interictal and preictal networks. Classification performances were assessed by AUC, the accuracy, and the F1-score and are presented in Tab. 1. In general, the connectivity patterns associated with activities of separated frequency bands cannot distinguish between preictal and interictal days. Only networks estimated from low $\gamma$ oscillations (30-49 Hz) performed better than a random classifier. Nevertheless, classification performances considerably increased when the information from different frequency bands was combined for each patient: All patients reached AUCs of $\geqslant 0.78$ (mean AUC $=0.93$, mean F1 score $=0.79$, and mean accuracy $=0.85$). These findings suggest that combining connectivity across different frequency bands provides valuable information for distinguishing preictal from interictal epochs. Notably, the hyperbolic embedding method outperformed a standard machine learning model, specifically a support vector machine (SVM). The SVM model yielded an average AUC of $0.79$, an F1 score of $0.78$, and a mean classification accuracy of $0.77$, demonstrating the efficacy of the proposed hyperbolic network embedding approach in enhancing seizure risk prediction.
@@ III.2 Forecasting of upcoming seizure(s)
Forecasting results were consistent with those of our previous study, in which we demonstrated that brain connectivity obtained from vigilance-controlled resting-state recordings could accurately predict the risk of upcoming seizures [18]. The results, presented in Table. 2, demonstrate the forecasting performance obtained for the distinct frequency bands when the prediction models for each day were trained with data from all preceding days. Combining information from different frequency bands yielded a good predictor with an average Brier score and BSS of $0.12$ [min. = 0.003, max. = 0.48] and $0.36$ [min. $=-0.43$, max. $=0.70$], respectively. In comparison, the SVM model achieved a mean Brier score of $0.13$. It is worthy of notice that connectivity networks from the $\beta$ and low $\gamma$ band also provided good predictors, in agreement with previous findings [18]. These results demonstrate that our prediction model outperformed other approaches, including a standard machine learning model, in forecasting the risk of upcoming seizure(s).
@@ IV Discussion
In contrast to existing models that depend on long-term EEG recordings to monitor preictal dynamics, our findings demonstrate that brief resting-state EEG segments can yield clinically relevant information regarding the preictal state on a day-to-day basis. It is worth noting that a direct comparison to traditional seizure prediction methods is not entirely feasible due to methodological differences, particularly in the duration of recordings utilized. However, compared to alternative models such as the SVM classifier [18] or a simple anteriority-based model, our approach exhibits superior performance. Additionally, most EEG-based methods for seizure prediction provide forecasts within minutes to, at most, an hour before a seizure event [14, 15]. In contrast, our method provides a single daily forecast, offering an extended time window for intervention or monitoring throughout the day.
@@
The application of hyperbolic geometry for embedding brain networks offers novel insights into the structural organization of functional connectivity. Hyperbolic space is inherently suited to represent hierarchical and complex network architectures, characteristics commonly observed in biological systems such as the human brain. In our model, hyperbolic distances between nodes encode both the strength and the hierarchical nature of inter-regional connections, enabling the detection of network alterations that may be overlooked by Euclidean or other geometric representations. Our findings indicate that this geometric framework yields a precise characterization of connectivity changes, including those associated with epileptogenic processes.
While other non-Euclidean methods, such as Riemannian embeddings of covariance matrices, have been extensively employed in EEG analysis (particularly within brain-computer interface applications), the use of hyperbolic geometry to model functional connectivity, especially in the context of epilepsy, remains relatively underexplored. This work represents, to our knowledge, the first application of such embeddings for seizure prediction, opening promising perspectives for future studies.
Despite the promising classification and prediction performance demonstrated by our model, several limitations should be acknowledged. First, due to the heterogeneity in electrode implantation sites across patients, the development of a generalized, patient-independent prediction model remains infeasible at this stage. Second, the vigilance-controlled resting-state iEEG recordings were obtained from hospitalized individuals undergoing presurgical evaluation. Although no significant associations between seizure occurrence and medication tapering were identified in the original dataset [18], the potential influence of time-dependent factors and treatment adjustments cannot be entirely excluded—an inherent limitation in such clinical environments. Consequently, the generalizability of our approach to out-of-hospital or real-world scenarios is limited, as these contexts involve greater long-term variability and more heterogeneous clinical profiles.
The inclusion of larger datasets with vigilance-controlled recordings at different times of day could facilitate evaluation of diurnal variations in seizure risk, thereby enhancing clinical relevance. Furthermore, integrating additional physiological signals—such as heart rate variability, spontaneous baroreflex sensitivity, and electrodermal activity—may further improve predictive performance.
To summarize, our findings indicate that hyperbolic embedding of brain networks captures discriminative features between interictal and preictal states, offering a promising framework for accurate daily seizure risk forecasting.